OpenStudy (samigupta8):

Let F(X)=2x^3-3/2x^2-1/8x+3/16 and x=y/a then the least positive integral value of a for which coefficients of f(y)=0 become integral is

OpenStudy (samigupta8):

\[f(x)=2x^3-\frac{ 3 }{ 2 }x^2-\frac{ 1 }{ 8 }x+\frac{ 3 }{ 16} and x=\frac{ y }{ a }\]

OpenStudy (rsadhvika):

x = y/a y = xa \(f(y) = 2(xa)^3 - 3/2 (xa)^2 - 1/8 x + 3/16 \)

OpenStudy (samigupta8):

don't u think there should be a even after

OpenStudy (samigupta8):

1/8x

OpenStudy (rsadhvika):

I am not sure f(y) = 0 in the question is throwing me off -- if you set f(y) to 0, you can multiply anything both sides and always have integral coefficients :o

OpenStudy (dumbcow):

f(y) needs to be in terms of y \[f(y) = \frac{2}{a^{3}}y^{3}-\frac{3}{2a^{2}}y^{2}-\frac{1}{8a}y+\frac{3}{16} = 0\]

OpenStudy (samigupta8):

y=xa then why r u not putting value as eqn multiplied by a and leaving that constant term

OpenStudy (dumbcow):

?? oh i got it, multiply equation by a^3 \[2y^{3}-\frac{3a}{2} y^{2}-\frac{a^{2}}{8}y+\frac{3a^{3}}{16}=0\] find integer whose cube is multiple of 16, square is multiple of 8, and even a = 4

OpenStudy (rsadhvika):

we can as well multiply the whole equation by 16 and get rid of fractions right ?

OpenStudy (dumbcow):

yes but they want a value for "a" to get rid of fractions

OpenStudy (rsadhvika):

asking for integral coefficients of f(y) = 0 makes no sense to me still

OpenStudy (rsadhvika):

question should be asking for integral coefficients of f(y) i think

OpenStudy (dumbcow):

@samigupta8 , do you see my posts above?

OpenStudy (samigupta8):

why u multiplied whole equation by a^3

OpenStudy (dumbcow):

so that "a" is in numerator ... impossible to make 2/a^3 an integer

OpenStudy (samigupta8):

okk...

OpenStudy (dumbcow):

sorry not impossible but a=1 wouldnt work for other coefficients its just a way to manipulate equation to get integer coefficients